Your family of 6 went to a Italian restaurant for dinner. Everyone had a sub, either meatball for $9 or chicken parmesan for $11. If the total bill was $58, how many people ordered each type of sub?

Answer by ReadingBoosters(3246)

9x+11y = 58
x+y=6
...
9(6-y) + 11y = 58
54 - 9y + 11y = 58
2y = 58 - 54
2y = 4
y = 2 people who ordered chicken parmesan
...
9x+11(6-x) = 58
9x + 66 - 11x = 58
-2x = 58 - 66
-2x = -8
x = 4 poeple whow ordered meatball
.....................
Delighted to help.
-Reading Boosters
Website: www.MyHomeworkAnswers.com
Wanting for others what we want for ourselves.